A=lw length=4cm width=3cm A=108cm squared what is the length and width? ratio is 3:4
Answer by checkley77(12844)

A=3N*4N
108=12N^2
N^2=108/12
N^2=9
N=SQRT9
N=3 THE MULTIPLIER FOR THE RATIOS.
3*3=9 ANS. FOR THE LENGTH.
4*3=12 ANS. FOR THE WIDTH.
PROOF:
108=9*12
108=108